# Federated Instruction Tuning of LLMs with Domain Coverage Augmentation

### v4

@@ Abstract
Federated Domain-specific Instruction Tuning (FedDIT) utilizes limited cross-client private data together with server-side public data for instruction augmentation, ultimately boosting model performance within specific domains. To date, the factors affecting FedDIT remain unclear, and existing instruction augmentation methods primarily focus on the centralized setting without considering distributed environments. Our experiments reveal that the cross-client domain coverage, rather than data heterogeneity, drives model performance in FedDIT. In response, we propose FedDCA, which optimizes domain coverage through greedy client center selection and retrieval-based augmentation. For client-side computational efficiency and system scalability, FedDCA, the variant of FedDCA, utilizes heterogeneous encoders with server-side feature alignment. Extensive experiments across four distinct domains (code, medical, financial, and mathematical) substantiate the effectiveness of both methods. Additionally, we investigate privacy preservation against memory extraction attacks utilizing various amounts of public data. Results show that there is no significant correlation between the volume of public data and the privacy-preserving capability. However, as the fine-tuning rounds increase, the risk of privacy leakage reduces or converges.
@@ 1 Introduction
Table 1: Performance(%) of different augmentation settings on each domain, conducted via FedAvg protocol with 10 clients. Specifically, the zero-shot directly inferences without any fine-tuning, while the base data method utilizes only the client’s local data for FedDIT. Additionally, we compare FedDCA with other two augmentation strategies: random sampling and direct retrieval (described in Appendix A.3), respectively.
Metric Zero-shot Base Data Random Sampling Direct Retrieval FedDCA
HumanEval/Pass@1 29.88 39.03 32.93 34.14 36.58
MMLU-Med/Acc 70.60 68.40 71.30 72.20 74.50
FPB/Acc 55.94 58.25 64.19 66.31 67.24
FiQA/Acc 18.54 14.18 13.09 19.11 35.27
TFNS/Acc 59.21 66.62 65.53 67.62 73.32
GSM8K/Exact Match 23.27 47.46 47.38 50.87 52.46
Avg. 42.91 48.99 49.07 51.71 56.56
@@
Recently, federated instruction tuning (FedIT) has gained attention as a novel approach that leverages the principles of federated learning (FL) to facilitate collaborative training of large language models (LLM) in distributed environments while maintaining the confidentiality of private data (McMahan et al., 2017; Ye et al., 2024b). This methodology allows for the exchange of model parameters among distributed data holders, thereby achieving a careful balance between privacy preservation and efficient model optimization. Despite the establishment of various FedIT frameworks (Ye et al., 2024b; Kuang et al., 2023; Zhang et al., 2023c), existing literature has not adequately addressed the practical challenges that Federated Domain-specific Instruction Tuning (FedDIT) may encounter in real-world applications. For instance, FedIT generally necessitates a sufficient amount of instruction data for fine-tuning, which is often a shortage in domain-specific fine-tuning contexts (Zhang et al., 2024b).
@@
In this study, we investigate FedDIT, a novel approach within the FL paradigm aimed at boosting the performance of LLMs in specific domains. Unlike general FedIT, which seeks to enhance model effectiveness across diverse tasks without accounting for local data shortage, FedDIT encounters the unique challenge of clients possessing only a limited quantity of local domain-specific data. To overcome this, FedDIT leverages a server-hosted public dataset that spans multiple domains to enrich the local data through a specific instruction augmentation strategy. This strategic enrichment is crucial for achieving effective instruction tuning and needs to be meticulously designed to avoid performance degradation. Except for the code domain, which primarily adheres to a standardized paradigm, our results reveal that when clients rely solely on their local data for FedDIT, even the presence of high-quality in-domain local data can be insufficient due to limited scale, leading to a decline in performance, as reflected in the underlined values in Table 1. In summary, the goal of FedDIT is to develop a domain-specific LLM that employs collaborative training and instruction augmentation with public data while safeguarding client privacy, thereby ensuring the model’s proficiency in executing tasks pertinent to its designated domain.
Additionally, the factors affecting FedDIT are still unclear. Compounding this uncertainty, introducing augmented instructions may further complicate results, making it difficult to ascertain effective improvement strategies. Shepherd (Zhang et al., 2023c) approaches this problem from the perspective of heterogeneity, constructing heterogeneity based on the topic of general instruction datasets. It demonstrates that, unlike the consensus of traditional FL, for general FedIT, heterogeneity has a positive effect. By aggregating diverse instructions from clients, the diversity increases, thereby enhancing the model’s adaptability to various tasks. However, it just scratches the surface and does not explore issues in FedDIT.
@@
Going one step further, we conduct experiments to unveil a significant finding (Appendix A.2): there is no linear correlation between the degree of non-independent and identically distributed (non-iid) and LLM’s performance in the context of FedDIT. Inspired by Explore-Instruct (Wan et al., 2023), which shows the potential of domain coverage in domain-specific instruction tuning. The cross-client domain coverage metric is initially defined, followed by an investigation into its impact on FedDIT. Results demonstrate that domain coverage significantly influences model performance in the corresponding domain.
@@
To maximize the cross-client domain coverage without compromising client data privacy, we propose a novel FedDIT algorithm, Federated Instruction Tuning of LLMs with Domain Coverage Augmentation, termed FedDCA. This algorithm employs a greedy client center selection process and implements instruction augmentation through dense retrieval on the server side. The fundamental idea of FedDCA is to select client centers to expand the diversity and coverage of augmented instruction datasets within a specific domain. By strategically optimizing domain coverage at each step, FedDCA efficiently constructs the augmented train set that enhances both the learning and generalization capabilities of the model, leading to superior performance on domain-specific tasks. Furthermore, to mitigate computational overhead on the client side and enhance the system scalability, we propose FedDCA, which employs heterogeneous encoders of different sizes and capacities. To achieve feature alignment, we train a projector on the server side using public data and employ contrastive learning techniques.
We demonstrate the effectiveness of FedDCA through comprehensive experiments conducted across four domains: code, medical, financial, and mathematical. These are compared against a range of baselines, which can be categorized into unaugmented and augmented methods. In the unaugmented setting, our method is compared with FedIT, which includes four orthodox FL techniques: FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), SCAFFOLD (Karimireddy et al., 2020), and FedAvgM (Hsu et al., 2019). In the augmented setting, we compare FedDCA against methods such as random sampling, direct retrieval, LESS (Xia et al., 2024), and Self-Instruction (Wang et al., 2022). Additionally, we present the performance outcomes of FedDCA when applied under various FL strategies. We also compare the computational efficiency on the client side between FedDCA and its variant, FedDCA. For privacy analysis, experiments against memory extraction attacks are conducted to evaluate how different quantities of retrieved public data affect the privacy of client local data. Results indicate that while reliance on local data increases memorization of sensitive information, the risk of privacy leakage diminishes or converges in the augmented setting as the training rounds progress.
The main contribution is as follows:
  • We reveal a critical finding: in the context of FedDIT, there exists no linear correlation between the degree of heterogeneity and model performance. Instead, cross-client domain coverage substantially impacts LLM’s effectiveness, as elaborated in Appendix A.2.
  • We propose a novel FedDIT algorithm (Section 4), termed FedDCA, aimed at maximizing cross-client domain coverage through greedy client center selection followed by retrieval-based instruction augmentation executed on the server. Additionally, we introduce FedDCA to further lessen the client-side computational overhead while enhancing the system scalability. This variant utilizes a heterogeneous encoder structure, paired with a projector on the server side for feature alignment.
  • Through extensive experiments (Section 5), we demonstrate the effectiveness of FedDCA and FedDCA. We also investigate privacy preservation against memory extraction attacks, conducting experiments based on various amounts of public data. Results suggest that the capacity for privacy preservation does not correlate significantly with the quantity of public data. In contrast, the risk of privacy leakage tends to decrease or converge as the fine-tuning rounds increase.
2 Preliminaries
FedDIT aims to leverage cross-client private domain-specific instruction data and utilize the multi-domain public data on the server to achieve instruction augmentation, collaboratively enhancing the model’s performance in specific domains. Consider $N$ distributed clients, each with local private data $D_{k}^{l}$. The server maintains a public dataset $D^{p}$ that encompasses multiple domains and is responsible for implementing data augmentation strategies and aggregating model parameters received from clients. The training process follows the standard FL protocol (McMahan et al., 2017). For computational efficiency, we adopt Low-Rank Adaption (LoRA) (Hu et al., 2022) as the fine-tuning method, which involves tuning additional parameters $\Delta\phi$ while keeping the pre-trained LLM’s parameters $\phi$ frozen. In the initial training phase, the server dispatches $\phi$ and $\Delta\phi$ to each client for $\mathcal{R}$ training rounds. In the $t$-th round, the server sends the aggregated $\Delta\phi^{t}$ to clients. Clients use $\Delta\phi^{t}$ to update their local LoRA parameters $\Delta\phi_{k}^{t}$ and conduct instruction tuning based on augmented instruction datasets $D_{k}$. Subsequently, the clients return $\Delta\phi_{k}^{t+1}$ to the server. The server then aggregates $\{\Delta\phi_{k}^{t+1}\mid k=1,\dots,N\}$ to obtain $\Delta\phi^{t+1}$ for the next round.
@@ 3 Problem Formulation
For better understanding, we list the frequently used notation in Appendix A.11. The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data (Ye et al., 2024b; Zhang et al., 2024b). Under the FL framework, suppose we have $N$ clients, where each client $c_{k}$ has a local dataset $D_{k}^{l}$ with its size $N_{k}^{l}$, and an augmented dataset $D_{k}^{g}$ from the server public dataset $D^{p}$, respectively. Due to constraints such as memory, computational overhead, and maximum tolerated training time, client $c_{k}$ can accept at most $N_{k}^{p}$ public instructions. Denote the augmentation strategy as $\Lambda$, through which the server performs instruction augmentation on the public dataset. If $\Lambda$ is null, it indicates that clients conduct FedDIT solely based on their local private data, which may lead to performance degradation. Conversely, $\Lambda$ may be classified into two categories: (1) focusing exclusively on the client’s own local data distribution or (2) considering the cross-client data distribution. In conclusion, the global objective of FedDIT is defined as follows:
$\operatorname*{arg\,min}_{\Delta\phi}{\left\{F(\phi,\Delta\phi)\triangleq\sum_{k=1}^{N}p_{k}\left((1-\alpha_{k})F_{k}\left(\phi,\Delta\phi_{k};D_{k}^{l}\right)+\alpha_{k}F_{k}\left(\phi,\Delta\phi_{k};D_{k}^{g}\right)\right)\right\}},$ (1)
where the first term and second term denote the accumulated fine-tuning loss computed on the client $c_{k}$’s local instructions $D_{k}^{l}$ and the augmented public instructions $D_{k}^{g}$ from the server, respectively. $p_{k}$ is the weight of the $k$-th client, which is determined by the ratio of $k$-th client’s data size to all clients’ total data size. $\alpha_{k}$ is the ratio of the public data amount of its augmented instruction dataset $D_{k}$, which is computed as $\frac{N_{k}^{p}}{N_{k}^{l}+N_{k}^{p}}$. Eq.1 minimizes the summed empirical loss across clients’ augmented instructions to pursue the in-domain utility of the obtained global model. Denote $P$ as the model parameters and $\mathcal{D}$ as a specific dataset, then the client $c_{k}$’s empirical loss $F_{k}(\phi,\Delta\phi_{k};\mathcal{D})$ base on $\mathcal{D}$ is calculated as:
@@
where $x_{j}\in D,\forall j\in\{1,2,\dots,|D|\}$. The instruction tuning loss $l(\cdot;\cdot)$ on a sample $(x,y)$ is defined as $-\sum_{t=1}^{\left|y\right|}\log\left(w\left(y_{t}|x,y_{<t}\right)\right)$, where $x$ is the formated instruction with Alpaca instruction template [^1] https://crfm.stanford.edu/2023/03/13/alpaca.html and $y$ is the corressponding response.
In contrast to the problem formulations in many prior works on FedIT, the primary distinction in the formulation of FedDIT lies in acknowledging the lack of in-domain instructions across clients. FedDIT establishes a public multi-domain dataset on the server for domain-specific instruction augmentation by a specific sampling strategy $\Lambda$. Therefore, each client only needs to hold a few high-quality domain-specific private data to collaborate with other clients to obtain a strong domain-specific LLM.
@@ 4 Method
We propose Federated Instruction Tuning of LLMs with Domain Coverage Augmentation (FedDCA), which enhances domain coverage to obtain a LLM that performs well on domain-specific tasks. FedDCA follows the standard FedIT protocol(Ye et al., 2024b) to perform federated instruction tuning. The novel design of FedDCA lies in the phase before training, which consists of two key modules: greedy client center selection and domain data retrieval, which are elaborated in the following; also see details in Algorithm 1 and Figure 1. For computational efficiency and system scalability, we introduce a variant of FedDCA with heterogeneous encoder setting, named FedDCA. We first formulate the optimization problem to solve for FedDCA.
Figure 1: Overview of FedDCA, which consists of three stages: 1) The client $c_{k}$ performs local instructions clustering and sends cluster centers $\mathcal{C}_{k}$ to the server. 2) The server first does greedy client center selection to maximize domain coverage and performs client-center-based domain data retrieval, then sends the augmented instructions $D_{k}^{g}$ to the client $c_{k}$. 3) Clients fine-tune the LLM collaboratively, and the LoRA parameters $\Delta\phi$ are exchanged between clients and the server.
Algorithm 1 FedDCA: greedy client center selection & domain data retrieval
1:Parameters: Number of clusters $\xi$; Encoder model $w_{enc}$; Client local datasets $D=\{D_{1},D_{2},\ldots,D_{N}\}$; Public dataset $D^{p}$; Similarity score threshold $\alpha$; Number of public data samples retrieved per client $N_{k}^{p}$; Client centers $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{N}\}$; Pre-encoded public instruction embeddings $\mathcal{E}$.
2:for $i\in\{1,2,\ldots,N\}$ do
3:     $\mathcal{E}^{\prime}\leftarrow\{w_{enc}(x)\mid x\in D_{i,instruction}^{l}\}$
4:     $\mathcal{C}_{i},S_{i}\leftarrow\text{k-means}(\xi).\text{fit}(\mathcal{E^{\prime}})$triangleright Cluster local instructions, return cluster centers $\mathcal{C}_{i}$ and sizes $S_{i}$
5:end for
6:Send $\{\mathcal{C}_{i},S_{i}\mid i\in\{1,2,\ldots,N\}\}$ to the server for the greedy client center selection
7:$\mathcal{P}\leftarrow\{\mathcal{C}_{0,k}\ \mid k=\arg\max(S_{0})\}$triangleright Initialize the client center set
8:for $i\in\{1,2,\ldots,N-1\}$ do $\triangleright$ Greedy client center selection
9:     $\mathcal{S}\leftarrow\sum(\mathcal{C}_{i}\cdot\mathcal{P}^{T},\text{dim}=-1)$triangleright Compute summed similarity score of each cluster center
10:     $\mathcal{I}\leftarrow(N-i)\text{-}\arg\text{sort}(\mathcal{S})$triangleright Filter $i$ cluster centers close to the client center set $\mathcal{P}$
11:     $\mathcal{I^{\prime}}\leftarrow\arg\text{sort}(-S_{i})$triangleright Sort cluster center by cluster size in descending order
12:     $j\leftarrow\text{first element of }\mathcal{I}\cap\mathcal{I}^{\prime}$triangleright Selected cluster center $\mathcal{C}_{i,j}$
13:     $\mathcal{P}\leftarrow\mathcal{P}\cup\mathcal{C}_{i,j}$triangleright Update the client center set
14:end for
15:for $i\in\{1,2,\ldots,N\}$ do $\triangleright$ Client center based domain data retrieval
16:     $\mathcal{S}\leftarrow\mathcal{P}_{i}\cdot\mathcal{E}^{T}$triangleright Compute similarity score between $\mathcal{P}_{i}$ and $\mathcal{E}$
17:     $\mathcal{S^{\prime}}\leftarrow\{s\mid s\in\mathcal{S},s<\alpha\}$triangleright Filter instructions with similarity score larger than $\alpha$
18:     indices of $\mathcal{I}\leftarrow\text{indices of }N_{k}^{p}\text{-top in }\mathcal{S^{\prime}}$
19:     $D_{i}^{g}\leftarrow\{D^{p}_{j}\mid j\in\mathcal{I}\}$
20:     $D_{i}\leftarrow D_{i}^{l}\cup D_{i}^{g}$triangleright Obtain the augmented instruction dataset $D_{i}$
21:end for

### v5

@@ Abstract
Federated Domain-specific Instruction Tuning (FedDIT) utilizes limited cross-client private data together with various strategies of instruction augmentation, ultimately boosting model performance within specific domains. To date, the factors affecting FedDIT remain unclear, and existing instruction augmentation methods primarily focus on the centralized setting without considering distributed environments. Our experiments reveal that the cross-client domain coverage, rather than data heterogeneity, drives model performance in FedDIT. In response, we propose FedDCA, which optimizes domain coverage through greedy client center selection and retrieval-based augmentation. At its core, the greedy selection procedure iteratively picks client centers that maximize the diversity and coverage of the instruction space while avoiding redundancy with previously selected centers. This ensures broad yet efficient coverage of the domain distribution across clients. For client-side computational efficiency and system scalability, FedDCA, the variant of FedDCA, utilizes heterogeneous encoders with server-side feature alignment. Extensive experiments across code, medical, financial, and mathematical domains substantiate the effectiveness of both methods, as well as plug-and-play capability. We further analyze privacy preservation against memory extraction attacks, showing that while privacy leakage risk is independent of augmented public data ratio, it decreases or converges as training progresses.
@@ 1 Introduction
Recently, federated instruction tuning (FedIT) has gained attention as a novel approach that leverages the principles of federated learning (FL) to facilitate collaborative training of large language models (LLM) in distributed environments while maintaining the confidentiality of private data McMahan et al. (2017); Ye et al. (2024b); Zhang et al. (2023b). This methodology allows for the exchange of model parameters among distributed data holders, thereby achieving a careful balance between privacy preservation and efficient model optimization. Despite the establishment of various FedIT frameworks Ye et al. (2024b); Kuang et al. (2023); Zhang et al. (2023b), existing literature has not adequately addressed the practical challenges that Federated Domain-specific Instruction Tuning (FedDIT) may encounter in real-world applications. For instance, FedIT generally necessitates a sufficient amount of instruction data for fine-tuning, which is often a shortage in domain-specific fine-tuning contexts Zhang et al. (2024b).
Figure 1: Visualization of different instruction augmentation strategies in the mathematical domain. The Random Sampling strategy randomly samples instructions from the public dataset. Direct Retrieval performs dense retrieval from the public dataset based on the client’s local data distribution without considering the cross-client domain coverage. FedDCA strategically selects client centers to maximize cross-client domain coverage and then performs domain data retrieval. The t-SNE visualization shows FedDCA has a better domain coverage than other sampling strategies.
@@
In this study, we investigate FedDIT, a novel approach within the FL paradigm aimed at boosting the performance of LLMs in specific domains. Unlike general FedIT, which enhances model effectiveness across diverse tasks without accounting for local data shortages, FedDIT addresses the unique challenge of clients only having limited domain-specific data. To overcome this, FedDIT enriches local data through a specific instruction augmentation strategy, crucial for effective instruction tuning and avoiding performance degradation. Our results reveal that, except for the code domain, where a standardized paradigm is primarily adhered to, relying solely on high-quality in-domain local data can be insufficient due to limited scale, leading to a decline in performance (see underlined values in Table 2). FedDIT aims to develop a domain-specific LLM through collaborative training and instruction augmentation while safeguarding client privacy, ensuring proficiency in tasks pertinent to its designated domain. For simplification and better instruction quality Zhang et al. (2024b); Toshniwal et al. (2024), we focus on a specific scenario of FedDIT involving a server-hosted public dataset, abstracting open-source instruction datasets on the web across multiple domains. This dataset is used for various instruction augmentation strategies, enhancing model performance in specific domains.
@@
Additionally, the factors affecting FedDIT are still unclear. Compounding this uncertainty, introducing augmented instructions may further complicate results, making it difficult to ascertain effective improvement strategies. We conduct experiments to unveil a significant finding: there is no monotonic correlation between the degree of non-independent and identically distributed (non-iid) and LLM’s performance in the context of FedDIT. Inspired by Explore-Instruct Wan et al. (2023), which highlights the potential of domain coverage in domain-specific instruction tuning. Unlike previous metric that measures domain coverage using the distribution of verb-noun pairs, we define a more general, representation-based cross-client domain coverage metric and investigate its impact on FedDIT. Our results show that domain coverage significantly influences model performance within the corresponding domain.
@@
To maximize the cross-client domain coverage without compromising client data privacy, we propose a novel FedDIT algorithm, Federated Instruction Tuning of LLMs with Domain Coverage Augmentation, termed FedDCA. This algorithm employs a greedy client center selection process and implements instruction augmentation through dense retrieval on the server side. The fundamental idea of FedDCA is to select client centers to expand the diversity and coverage of augmented instruction datasets within a specific domain (as shown in Figure 1). By strategically optimizing domain coverage at each step, FedDCA efficiently constructs the augmented train set that enhances both the learning and generalization capabilities of the model, leading to superior performance on domain-specific tasks. Furthermore, to mitigate computational overhead on the client side and enhance the system scalability, we propose FedDCA, which employs heterogeneous encoders of different sizes and capacities. To achieve feature alignment, we train a projector on the server side using public data and employ contrastive learning techniques.
We demonstrate the effectiveness and plug-and-play capability of FedDCA through comprehensive experiments conducted across four domains: code, medical, financial, and mathematical. These are compared against a range of baselines, which can be categorized into unaugmented and augmented methods. Results show that FedDCA outperforms other methods from at least 0.84% to the maximum of 29.19% in terms of model performance, with a relative domain coverage improvement from 4.82% to 21.36%. Extensive experiments across privacy analysis, computational efficiency and other further investigations are also conducted.
@@ 2 Problem Formulation
Federated Domain-specific Instruction Tuning (FedDIT) is a federated learning approach designed to improve the performance of LLMs in specific domains by utilizing limited cross-client private data in combination with domain-specific instruction augmentation strategies Zhang et al. (2024b); Xia et al. (2024); Wang et al. (2022). We abstract the public data as a server-hosted multi-domain dataset (the composition is detailed in Table 1). The presence of a public dataset on the server is only one possible scenario, and the proposed algorithm is independent of the presence of a public dataset on the server. Even if the server does not have a public dataset, clients can retrieve public instructions by themselves based on the received client centers from the server or perform self-instruct, thereby achieving data augmentation.
Consider $N$ distributed clients, each with local private data $D_{k}^{l}$ of size $N_{k}^{l}$, and augmented data $D_{k}^{g}$ from the server’s public dataset $D^{p}$. Due to constraints like memory and computation, client $c_{k}$ can accept at most $N_{k}^{p}$ public instructions. The server maintains $D^{p}$ that spans multiple domains and is responsible for data augmentation strategies $\Lambda$ and parameter aggregation. For efficiency, we adopt Low-Rank Adaption (LoRA) Hu et al. (2022) as the fine-tuning method, tuning additional parameters $\Delta\phi$ while keeping pre-trained LLM parameters $\phi$ frozen.
The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data Ye et al. (2024b); Zhang et al. (2024b, 2023b), which is defined as:
$\small\operatorname*{arg\,min}_{\Delta\phi}\left\{F(\phi,\Delta\phi)\triangleq\sum_{k=1}^{N}p_{k}F_{k}\left(\phi,\Delta\phi_{k};D_{k}^{l},D_{k}^{g};\alpha_{k}\right)\right\},$ (1)
where $F_{k}(\cdot)$ is the accumulated instruction fine-tuning loss of model $P_{\phi+\Delta\phi_{k}}$ computed on the client $c_{k}$’s augmented dataset $D_{k}^{l}\cup D_{k}^{g}$. $p_{k}$ is client $k$’s weight based on data ratio, and $\alpha_{k}=\frac{N_{k}^{p}}{N_{k}^{l}+N_{k}^{p}}$ is the ratio of public data. Eq.1 minimizes the summed empirical loss across clients’ augmented instructions to pursue the in-domain utility of the obtained global model. The client $c_{k}$’s empirical loss $F_{k}(\phi,\Delta\phi_{k};\mathcal{D})$ is calculated as:
@@
where $x_{j}\in D,\forall j\in\{1,2,\dots,|D|\}$. The instruction tuning loss $l(\cdot;\cdot)$ on a sample $(x,y)$ is defined as $-\sum_{t=1}^{\left|y\right|}\log\left(w\left(y_{t}|x,y_{<t}\right)\right)$, where $x$ is the formated instruction and $y$ is the corressponding response.
In summary, we formalize FedDIT as a novel FL framework that focuses specifically on domain-specific instruction tuning. In contrast, FedIT typically demands a substantial quantity of instructions for fine-tuning. Domain-specific instruction augmentation can be divided into two categories: 1) Generative methods, which generate instructions through the pre-trained LLM locally or external APIs. 2) Retrieval-based methods, which retrieve instructions from the web. The former methods either compromise privacy or have high computational overhead. Thus, in this work, we focus on retrieval-based methods to utilize diverse and high-quality public data.
3 What Truly Counts in FedDIT
(a) Exp1. Code
(b) Exp1. Med
(c) Exp1. Fin
(d) Exp1. Math
(e) Exp2. Performance (%)
(f) Exp2. Domain Coverage
Figure 2: (a)-(d) Performance (%) of different heterogeneity in each domain with 10 and 100 clients. (e)-(f) Performance and domain coverage of iid and non-iid settings on different domains. Experiments show that cross-client domain coverage significantly impacts model effectiveness, while data heterogeneity shows no monotonic correlation with performance.
Data Heterogeneity Is Not Matter in FedDIT.
We adopt the Dirichlet distribution Wang et al. (2020); Yurochkin et al. (2019) to construct various heterogeneity ($\alpha=[0.01,0.1,1,10]$) and use k-means with $\xi=100$ to pseudo labeling instructions.
We perform instruction tuning on both 10 and 100 clients with 2 randomly selected clients participating in each round. For each domain, we only use the in-domain data and then perform FedIT. We perform the experiments 3 times with different random seeds (42, 43 and 44) and report the average performance and the standard deviation in each domain. As shown in Figures 2(a), 2(b), 2(c) and 2(d), the performance of LLM does not decrease due to the increase of data heterogeneity but shows a non-monotonic correlation, which indicates that the performance of LLM does not directly depend on data heterogeneity and other factors that play a key role.
Domain Coverage: A Key Factor in FedDIT.
Different from Explore-Instruct, which defines domain coverage through the distribution of verb-noun pairs, we attempt to conduct more in-depth and extensive experiments to study the effect of domain coverage on FedDIT. Firstly, we define the domain coverage in the FL setting, considering the cross-client data distribution. Assume the dataset of in-domain data $D^{d}$ represents the latent data distribution of this domain and the cross-client data is defined as $D^{c}=\cup_{k=1}^{N}\left(D_{k}^{l}\cup D_{k}^{g}\right)$.
Inspired by the facility location function Cornuejols et al. (1983), we define the domain coverage of $D^{c}$ respect to $D^{d}$ as follows:
$\small d(D^{d},D^{c})=\frac{1}{\left|D^{d}\right|}\sum_{d\in D^{d}}\text{max}_{v\in(D^{c}\cap D^{d})}sim(d,v),$ (3)
where $sim(d,v)$ is the cosine similarity between $d$ and $v$. Note that we only use the in-domain data in $D^{c}$ for the domain coverage calculation to avoid misleading.
To better align with the real-world scenarios, we explore instruction augmentation based on both iid and non-iid cross-client data distribution and then perform dense retrieval for FedDIT. We set the number of clients to 10, while each client has 100 local instructions and obtains 1000 augmented public instructions from the server. For the iid setting, we randomly sample 1k from the public dataset and divide them into 10 shards as each client’s local data. For the non-iid distribution, we perform k-means clustering with $\xi=100$. Each client randomly samples 100 instructions from randomly selected distinct clusters.
Figures 2(e) and 2(f) presents the performance and the according domain coverage of FedDIT in different domains with iid and non-iid settings. We can observe that both iid and non-iid settings outperform in some domains, but both collectively indicate that higher domain coverage correlates with better performance.
@@ 4 Method
Figure 3: Overview of FedDCA, which consists of three stages: 1) The client $c_{k}$ performs local instructions clustering and sends cluster centers $\mathcal{C}_{k}$ to the server. 2) The server first does greedy client center selection to maximize domain coverage (obtains $\mathcal{P}$) and performs client-center-based domain data retrieval, then sends the augmented instructions $D_{k}^{g}$ to the client $c_{k}$. 3) Clients fine-tune the LLM collaboratively, and the LoRA parameters $\Delta\phi$ are exchanged between clients and the server.
Based on the above empirical observations, we propose FedDCA, which enhances domain coverage to obtain a LLM that performs well on domain-specific tasks (shown in Figure 3). We first formulate and analyze the cross-client domain coverage optimization problem. Then, we introduce the FedDCA algorithm, which consists of two main stages: greedy client center selection and client-center-based domain data retrieval. For computational efficiency and system scalability, we introduce a variant of FedDCA with heterogeneous encoder settings, named FedDCA.